### v2

# Modeling the Solar System as an Observed Multi-Transit System I: Characterization Limits from Analytic Timing Variations

Abstract
Planetary systems with multiple transiting planets are beneficial for understanding planet occurrence rates and system architectures. Although we have yet to find a solar system (SS) analog, future surveys may detect multiple terrestrial planets transiting a Sun-like star. In this work, we simulate transit timing observations of our system based on the actual orbital motions of Venus and the Earth-Moon Barycenter (EMB) — influenced by the other SS bodies — with varying noise levels and observing durations. We then retrieve the system’s dynamical parameters using an approximate N-body model for transit timing shifts while considering four possible plane-parallel configurations: 2-, 3-, 4-, and 5-planets.
[…]
Astrobiology (74), Exoplanet detection methods (489), Solar system planets (1260)
software: CalcEph(Gastineau et al., 2015), TTVFaster (Agol & Deck, 2016), Matplotlib (Hunter, 2007; Caswell et al., 2019), Julia (Bezanson et al., 2017)
1 Introduction
Early theories of planetary system formation and evolution were based exclusively on solar system (SS) observations (e.g. Lissauer & de Pater, 2013; Horner et al., 2020; Kane et al., 2021, and references therein). As our only reference at the time, astronomers expected extra-solar planetary systems to closely resemble our own: terrestrial planets in the inner regions, accompanied by natural satellites, as well as wide separation gas giants (with semi-major axes $>$3 au). Instead, the simultaneous survival of outer gas giants with multiple inner terrestrial companions is rarely seen in exo-planetary systems (e.g. Kepler-167 Chachan et al., 2022), possibly due to selection effects. In greater numbers we have discovered exoplanetary systems that display planets and planetary system architectures which differ dramatically from the SS; most notably: short-period gaseous planets (e.g. Mayor & Queloz, 1995; Wright et al., 2012), and a mix of rocky and close-in super-Earths and mini-Neptunes (e.g. Barnes et al., 2009; Howard, 2013).
The presence of gaseous planets in close-in orbits – where they could not have formed – resulted in a thorough re-examination of planet formation and evolution pathways (e.g. Adibekyan et al., 2013; Mordasini et al., 2015). Lin et al. (1996) introduced planetary migration to explain such objects, prompting astronomers to pay attention to the interactions of planetesimal cores (and planets) with each other and with the protoplanetary disk (Winn & Fabrycky, 2015). Astronomers have since applied such theories to our SS. By studying the impact of giant planet migration on terrestrial planet formation, as well as terrestrial survival during a system’s evolution (Mandell & Sigurdsson, 2003; Raymond et al., 2006; Mandell et al., 2007), we have hoped to reconcile the observed exoplanet population with the SS.
A cornerstone of these formation theories is the core-accretion theory for the formation and evolution of gas giant planets, both in our SS and beyond (Pollack et al., 1996). In this theory, Jupiter and Saturn formed by gravitationally capturing material from the protoplanetary disk onto their silicate/ice cores. Also, planetesimal-driven migration and dynamical instability are thought to have been responsible for their present-day orbits (see Raymond, 2024, and references therein). However, as we have yet to identify all the physical processes involved, it can be difficult to make the simulated systems match the observed outcomes (Raymond & Morbidelli, 2022). This motivates an empirical approach to the problem, so in this work we explore the simultaneous detection and characterization of a gas giant and a terrestrial planets within a single system using existing methods of planet discovery.
The vast majority of currently catalogued exoplanets have been detected using either the radial velocity (RV) Doppler shift or the primary transit signal. Here we focus on the transit technique, which was brought to the fore with the launch of NASA’s Kepler Space Telescope (Borucki et al., 2010). The intent of Kepler was to discover numerous planets whose orbits fall within the habitable zone (HZ) – the region around a star where a planet can sustain surface liquid water given sufficient greenhouse warming (Kasting, 1993). Although Kepler discovered some planets larger than Earth in the HZ of Sun-like (F/G/K) stars, it did not detect a true Earth-analogue planet, and likely neither will its successor, the Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015), which was designed to detect hotter super-Earths around bright, nearby stars. Thus, it remains to future, more sensitive surveys – either using transits (e.g., PLATO, Rauer et al. 2014; Nautilus, Apai et al. 2019) or direct-imaging (e.g., HabEx, Gaudi et al. 2019; LUVOIR, Roberge et al. 2021) – to find an Earth-Sun twin.
Additionally, while the techniques mentioned above are the most efficient for discovering exoplanets, our ability to characterize these planets is also observationally-limited. From the depth of a transit, we have measured the planets’ radii relative to the host star (Borucki et al., 2012; Quintana et al., 2014). However, we cannot measure masses for most single exoplanets that transit Kepler stars as they are too faint for RV follow-up observations. Although RV surveys have found Earth-mass planets orbiting bright, nearby M dwarf stars within their host-stars’ HZs (Mayor et al., 2014), the small probability of transit for such systems means that it is unlikely for astronomers to measure both the masses and radii of these planets. Therefore, our precise knowledge of the densities of individual planets is restricted by these techniques due to the reliance on the combining multiple observation types. Excluding the TRAPPIST-1 system and LHS 1140b (Agol et al., 2021; Lillo-Box et al., 2020), no exoplanets that reside in their host stars’ HZs have both measured masses and radii. Their bulk densities, atmospheric thicknesses, escape velocities, and surface gravities are thus unconstrained, which are all critical for assessing habitability. Sun-like stars have less activity and a shorter pre-main sequence than M-dwarf hosts like TRAPPIST-1 and LHS-1140, hence we would still like to find and characterize solar-system analogs.
In spite of the limited ability for inferring masses of most transiting exoplanets, an opportunity exists in systems with at least two transiting exoplanets, wherein gravitationally-induced planet-planet interactions can significantly perturb the transit times (e.g., Kepler-9 Ragozzine & Holman, 2019). The resulting transit timing variations (TTVs) were first recognized independently by Agol et al. (2005) and Holman & Murray (2005) as holding the promise of measuring masses of terrestrial exoplanets in a multi-transiting system (MTS) with near-resonant planets. The same TTV analysis can also reveal previously undetected planets, as shown initially by Ballard et al. (2011) and Nesvorný et al. (2012), and more recently by Heller et al. (2020). Indeed, this is the technique used to measure the masses of the TRAPPIST-1 planets.
Due to their sizes and proximity to their host, Venus and Earth are the most likely SS objects to be observed transiting at a distance (Wells et al., 2017). In this case, Venus and Earth would be among the longest period MTS planets. Therefore, if transits of similar planets were observed, could we use their TTVs to constrain their masses and orbits, as well as those of unseen companions? This work could inform future surveys – which push towards smaller masses/radii and longer orbital periods – and thereby characterize a more diverse group of systems than we have already found.
In this paper series, we will analyze the orbital perturbations induced by two transiting terrestrial planets on one another, Venus and the Earth – along with perturbations by SS objects that are otherwise undetectable – by treating the SS as an extra-solar system. We then test our ability to identify the presence and retrieve the parameters of these additional bodies, while simultaneously measuring the masses of the transiting planets. If we aim to search for SS-like systems, what better way to demonstrate how to characterize habitable planets around sun-like stars than by using our own SS?
For the purposes of this current paper, we model the Solar System (SS) as seen from afar, with one modification: we treat the Earth-Moon dynamical barycenter (EMB) as the physical location of the second observed planet. Our study of the Earth’s transit times as perturbed by the Moon will be investigated in a future paper, but for now we will assume that the EMB transit time can be measured from detecting transits of both the Earth and Moon, and fitting for the center of mass between them. In section 2.1, we describe our simulated transits of Venus and the EMB. Our description of the transit timing analysis in §2.2 includes summaries of the analytic model in § 2.2.2, and our likelihood estimate in § 2.2.3. We provide results from our retrieval in section 3, and discuss our findings and provide conclusions in section 4. We aim to find the required timing precision and observing schedule (i.e. survey design) to address the following questions:
  1. 1.
    Which non-transiting objects can we detect (e.g. Mars, Jupiter) and to what degree can we characterize them?
  2. 2.
    How precisely can we characterize the masses of the two transiting terrestrial planets, Venus and the EMB, amidst the presence of the other planets?
2 Methods
We simulate transit times from the dynamical orbits of both planets that would be viewed by a distant observer – assuming that there is a high signal-to-noise ratio for the stellar flux measurements. Using both the JPL ephemeris data and the CalcEph package, we compute the ephemerides for Venus and EMB relative to the Sun (Gastineau et al., 2015). Mercury and the remaining SS objects are included in the JPL ephemeris (see Fig. 1) and perturb Venus and the EMB. For each simulation that we run, our code does three things: checks for transits, measures the transit times, and adds some scaled Gaussian noise. We carry out simulations for a variety of survey durations and timing precision levels to explore how these impact our transit timing analysis. Starting with the longest observing duration and the smallest timing uncertainty, we progressively decrease our time-span along with varying the timing uncertainty.
We then retrieve the masses and orbital elements of the SS bodies by modeling the transit times with TTVFaster (Agol & Deck, 2016, hereafter AD16), and then carrying out fits to the transit times for a global search for additional planets using the Levenberg-Marquardt algorithm (LM Levenberg, 1944; Marquardt, 1963), followed by an affine-invariant approach of Markov Chain Monte Carlo (MCMC) analysis to characterize the uncertainties (Goodman & Weare, 2010). Assuming that an observer would be unaware of the existence of non-transiting objects, we fit the TTVs with a sequence of multi-planet system models. Beginning with a preliminary 2-planet model, we progressively add objects to the model – namely, Jupiter and Mars analogs. This step-by-step fitting allows us to analyze the SS as if it were an extra-solar MTS. Note that we assume the planets are plane-parallel and edge-on in their orbits, which is justified for our SS in which the RMS inclination of our eight planets is less than 2 (Winn & Fabrycky, 2015).
For each scenario, we perform a likelihood profile analysis to search for non-transiting objects, from which we estimate the maximum likelihood with respect to the parameter set. Subsequently, we carry out an affine invariant MCMC analysis to infer the model parameter probability distribution. Using both the likelihood profile and the MCMC sampling is beneficial for multiple reasons: the former checks for the modality of the probability distribution, and the latter robustly characterizes the model parameter uncertainties. In the following sections, we describe our methods in more detail.
2.1 Simulating data
2.1.1 Simulation specifications
The JPL ephemerides are generated by fitting numerically integrated orbits to both ground-based and space-based observations of SS bodies. For this paper, we adopt the DE440 version of the ephemerides, and employ CalcEph to compute planetary position, velocity, and acceleration vectors at a given Julian Date (JD; Park et al., 2021). A randomly-oriented, distant observer is unable to see all of the planets in the system transiting the Sun (Brakensiek & Ragozzine, 2016), so we calculate the line of sight required to observe the transits of both planets in appendix A.1, along which distant observers would see our SS like we see MTS.
Figure 1 shows our simulated system integrated for 15 yrs. Over numerous orbits, the location from which both Venus and the EMB can be seen transiting the Sun – dubbed the transit zone – would change very slightly due to planet-moon interactions, and planet-planet interactions (Heller & Pudritz, 2016). However, the individual transit zones for terrestrial SS planets would be valid for thousands of years (Wells et al., 2017).
To perform our simulations, we provide a length of time for which we want observations ($n_{\text{year}}$), starting at some time ($jd_{0}$). While the former varies, we fix the start time to $jd_{0}=$ 2433282.5 JD to maintain consistency across our analyses  [^1] The DE440 JPL ephemerides are accurate over the epochs JD 2287184.5-2688976.5.. We compute initial orbits from $jd_{0}$ stopping at $jd_{\text{end}}=365.25\times n_{\text{year}}+jd_{0}$ with a time step of 22.5 days. In this paper, our observation time-spans range from 15 to 30 years, which encompass at least one full orbit of Jupiter. This is a much longer observing duration compared to the periods of known exoplanets, but this length of time is needed to observe long-term TTVs, which may be applicable to systems within the TESS continuous viewing zone.
Figure 1: Top-down view of solar system orbital paths computed from JPL ephemerides over 15 yrs. Right panel: Observer line of sight required for planetary transits of Venus (triangles) and EMB (circles), calculated in Appendix A.1.
2.1.2 Mid-transit times
During the orbit integration described above, we track the Cartesian coordinates and velocities to compute the sky-projected planet-star separation vector, $\textbf{r}_{\text{sky,}i}$, and its time derivatives for the $i$th planet. Let us introduce the function $f=\textbf{r}_{\text{sky,}i}\cdot\dot{\textbf{r}}_{\text{sky,}i}$, which we derive in Appendix A.2. For each body (i.e. Venus and EMB), the first guess at the time of transit is the minimum of the derivative of $f$ within one planetary period of the start time $jd_{0}$ – dubbed $f^{\prime}$. Once a transit has been identified as having occurred between two orbit time steps, we refine the planet’s transit times by seeking the roots of $f$ with the Newton-Raphson method (see Kong et al., 2021). In brief, we start with the minimum of $f^{\prime}(t)$ as an estimate ($t=t_{0}$), and improve the root approximation by evaluating
$t_{n+1}=t_{n}-\frac{f(t_{n})}{f^{\prime}(t_{n})}\text{.}$ (1)
We iterate until the current time estimate, $t_{n+1}$, is equal to either of the two most recent estimates, $t_{n}$ or $t_{n-1}$.
After the first transit time is found, we advance the search time frame by one orbit ($P_{i}$), to find the next transit time. More transit times are found by adopting new start and end times until we reach the end of our ‘observation’ time, $jd_{\text{end}}$. For the $j$th transit observation of the $i$th planet, we collect the measured transit times from $jd_{0}$ and $jd_{\text{end}}$ for each planet as $t_{\text{obs},ij}$. Figure 1 shows the locations where Venus and the EMB could be simultaneously observed at the mid-transit point during their orbits about the Sun. The observer’s line of sight is represented by the dashed line.
2.1.3 Injected timing noise
Photometric noise will affect the accuracy to which an observer can measure the planetary mid-transit times, and consequently, the accuracy to which they can characterize the system. By simulating the effects of stellar brightness variations on the transits of an Earth-like planet across a Sun-like star, Morris et al. (2020) find that the PLATO mission would have an approximate 86-second [^2] Adopted from Figure A1 of Morris et al. (2020), which includes posterior distributions following a fit to the Sun’s stellar variability. error on Earth’s mid-transit time. This noise floor is mostly due to $p$-mode oscillations affecting the disc-integrated stellar flux used in modeling the light-curves. However, simulations show that the precision of mid-transit times can be improved by conducting multi-wavelength observations and decorrelating correlated noise across wavelength (e.g. Gordon et al., 2020; Borsato et al., 2022).
As such, we set our Gaussian uncertainties at 10, 30, 60, 90, and 120 seconds: covering values that range from optimistic to realistic. For each mid-transit time, we generate random normal deviates scaled by the selected timing uncertainty, which we then add to the synthetic transit times for the transit-timing analysis. Note that we use the same timing precision for both planets, $\sigma_{\text{obs}}$, as, in practice, they will differ only slightly. Although Venus’ transit is shallower relative to Earth, its faster orbital speed causes a sharper transit ingress and egress; this would yield timing precision similar to that of Earth.
Along with the number of transits for each planet, we summarize the simulations generated in Table 1. The total number of observations is $N_{\text{obs}}=\sum_{i=1}^{N_{p}}T_{i}$, where $T_{i}$ is the number of transits for planet $i$ and $N_{p}$ is the number of known planets – the subscripts $i=1$ and $i=2$ refer to Venus and the EMB, respectively. Each dataset is given by $\textbf{y}=(\left\{t_{\text{obs},ij},\sigma_{\text{obs}};j=1,...,T_{i}\right\};i=1,...,N_{p})$.
Table 1: Summary of solar system simulations.
$n_{\text{year}}$ $T_{1}$ $T_{2}$ $N_{\text{obs}}$
15 25 16 41
20 33 21 54
25 41 26 67
30 49 31 80
$\sigma_{\text{obs}}$ (sec) 10 30 60 90 120
  • Notes. With respect to the observing baseline, $T_{1}$ and $T_{2}$ are the number of transit time observations for each planet, respectively; $N_{\text{obs}}=T_{1}+T_{2}$ is the total number of observations. The last row shows the injected Gaussian noise level.
We compute TTVs by comparing the $t_{\text{obs},ij}$ to calculated transit times assuming a linear ephemeris, which we describe next.
2.2 Fitting transit times
2.2.1 Transit ephemeris
If the zeroth order Keplerian solution was correct for each planet, it would satisfy an ephemeris with a constant transit period, $P_{i}$. From this solution, one can calculate mid-transit times as
$t_{\text{calc},ij}=t_{0,i}+P_{i}\times E_{j},$ (2)
where $t_{0,i}$ is the initial time of the transit at epoch zero ($j=1$). The epoch, $E_{j}$, is an integer transit number equal to the number of orbits after the initial transit time for the $j$th transit. For each planet, we perform a linear regression of the observed transit times as a function of the epoch to derive an initial estimate for the mean ephemeris ($t_{0,i}$, $P_{i}$). Assuming this linear ephemeris, we produce calculated transit times for each planet, then compute observed minus calculated residuals for each transit time ($O-C$), i.e. the TTVs.
Table 2 shows expected transit times of the linear ephemeris, and deviations from the fixed period model for Venus and EMB. This simulation set was generated for a 30-yr timescale with Gaussian uncertainties at the 30-s level. In practice, transit-timing data often have missing or discontinuous observations, while in this paper we simulate contiguous transits based on the computed orbits. The TTV peak amplitudes observed – defined as the maximum of the absolute values– are $4.128$ min for Venus and $5.243$ min for EMB.
Table 2: Transit timing data for Venus (planet b) and the EMB (planet c).
Planet Transit Number $t_{\text{calc}}$ TTV Noise added to $t_{\text{obs},ij}$ $\sigma_{\text{obs}}$
[JD $-2430000$] [min] [min] [min]
Venus 0 3503.765486 -1.537 -0.28 0.5
Venus 1 3728.466261 -0.7195 -0.22 0.5
Venus 2 3953.167035 2.332 0.014 0.5
Venus 3 4177.867809 -0.579 -0.15 0.5
Venus 4 4402.568584 -2.802 0.89 0.5
Venus 5 4627.269358 2.964 -0.57 0.5
Venus 6 4851.970132 3.307 -0.23 0.5
Venus 7 5076.670907 1.332 0.078 0.5
Venus 8 5301.371681 -2.652 -1.3 0.5
Venus 9 $5$3.736 0.5 0.5
$\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$
EMB 0 3624.402641 3.928 -0.68 0.5
EMB 1 3989.659068 0.5787 0.15 0.5
EMB 2 4354.915494 1.524 0.61 0.5
EMB 3 4720.171921 -2.248 -0.23 0.5
EMB 4 5085.428348 1.259 -0.18 0.5
EMB 5 5450.684774 2.335 -0.29 0.5
EMB 6 5815.941201 -3.939 0.45 0.5
EMB 7 6181.197628 -3.393 0.12 0.5
EMB 8 6546.454054 -2.799 -0.098 0.5
EMB 9 6911.710481 -2.644 0.33 0.5
$\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$
  • Notes. Transit times were computed with CalcEph from JPL ephemerides over 30 yrs of solar system simulations, starting at $2433282.5$ JD. TTVs are found by linearly fitting the observed times with the injected Gaussian noise and finding the mean transit ephemeris. The entire table is available electronically in machine-readable format.
2.2.2 TTVFaster model
Next, we apply a retrieval to the simulated transit times, by using a dynamical model to infer the properties of the planetary system. The linear ephemeris from Eq. 2 is added to the perturbations on the transit times, the latter of which are computed from formulae derived in AD16. The TTVFaster package implements these analytic formulae given a set of planetary parameters per planet: the planet-to-star mass ratio ($\mu_{i}=m_{i}/M_{\star}$), the orbital period ($P_{i}$) in days, the initial time of transit ($t_{0,i}$), the eccentricity ($e_{i}$), and the longitude of periastron ($\varpi_{i}$) as measured from line of sight. In our edge-on and plane-parallel model, the argument of periastron ($\omega$) is equal to the longitude of periastron ($\varpi$); we adopt the $\omega$ variable going forward.
Assuming that the pairs of planets within a system are co-planar, the planets would each experience perturbations to their orbital frequencies, which correlate to perturbations in the times of transit. In this framework, the transiting planets are each perturbed by the other planets, whose orbits are held fixed, and which we approximate as epicycles. These are accurate assumptions for our SS as the timescales of our simulations are short compared with secular timescales, the eccentricities are small, and each pair of planets in our model is exterior to first-order mean-motion resonance.
Suppose that we observe two planets transiting their stellar host, and we measure the times of transit. Let the functions $F_{1,2}$ and $F_{2,1}$ represent the interactions between these two planets, where we label the inner planet as $i=1$, and the outer planet as $i=2$. AD16 defines the unperturbed orbital frequencies by $n_{i}^{2}=GM_{\star}/a_{i}^{3}$, where $n_{i}=2\pi/P_{i}$ are the mean motions. In addition to producing TTVs, the presence of another planet will also change the average orbital period of each planet. As such, we use the $\alpha=a_{1}/a_{2}\approx(P_{1}/P_{2})^{2/3}$ approximation when computing the TTVs for the pair of planets.
For each planet, we use the observed transit times to compute several values from AD16: initial longitudes, mean longitudes ($\lambda_{i}$), Laplace coefficients, and disturbing function coefficients. Assuming that the orbital reference is along the line of sight, we can define the mean longitude to first order in eccentricity by
$\lambda_{i}(t)=2\pi\Big{(}\frac{t-t_{0,i}}{P_{i}}\Big{)}+2e_{i}\sin{\omega_{i}},$ (3)
at a time $t$, and compute the difference in mean longitudes (i.e. $\psi=\lambda_{1}-\lambda_{2}$). In a two-planet system, the solution for the inner planet’s transit times is given by
$\delta t_{1,j}=\frac{P_{1}}{2\pi}\mu_{2}F_{1,2}(\alpha,\lambda_{1},\omega_{1},e_{1},\lambda_{2},\omega_{2},e_{2}),$
where both $\lambda_{1}$ and $\lambda_{2}$ are evaluated at the inner planet’s calculated transit times. Similarly, the outer planet’s times of transit are shifted by
$\delta t_{2,j}=\frac{P_{2}}{2\pi}\mu_{1}F_{2,1}(\alpha,\lambda_{2,},\omega_{2},e_{2},\lambda_{1},\omega_{1},e_{1}),$
where both $\lambda_{2}$ and $\lambda_{1}$ are evaluated at the outer planet’s calculated transit times. Note that we have defined the timing perturbations as $F_{1,2}$ and $F_{2,1}$ in units of the mass-ratio of the companion planet and the inverse orbital frequency of the transiting planet. For a detailed derivation of the equations above, please refer to AD16.
By summing over the perturbations for each pair of planets in a system of more than two planets, we can expand the above solution to approximate a multi-planet system. To describe the perturbations of some planet $k$ on the $i$th planet, we generalize $\alpha$ as the inner-outer planet period ratio such that $\alpha_{ik}=(P_{i}/P_{k})^{2/3}$ for $i<k$, and $\alpha_{ik}=(P_{k}/P_{i})^{2/3}$ for $i>k$. Thus, the TTVs for the $j$th transit of planet $i$ in a system with $N_{p}$ planets becomes
$\delta t_{ij}=\frac{P_{i}}{2\pi}\sum_{k\neq i}^{N_{p}}\mu_{k}F_{i,k}(\alpha_{ik},\lambda_{i},\omega_{i},e_{i},\lambda_{k},\omega_{k},e_{k}),$ (4)
where both longitudes ($\lambda_{i}$ and $\lambda_{k}$) are evaluated at the calculated transit times for the $i$th planet via Eq. 3. Afterwards, we add this solution back to the mean transit ephemeris to find model transit times for the $j$th transit of the $i$th planet,
$t_{\text{mod},ij}=t_{\text{calc},ij}+\delta t_{ij}\text{.}$
Note that we compute $\lambda$, $\alpha$, and $t_{\text{calc},ij}$ by allowing $t_{0,i}$ and $P_{i}$ to vary when optimizing our model.
In summary, the analytic formula used to compute a planet’s TTVs scales in proportion to the orbital period, the companion planet’s mass (relative to the star’s), both planets’ eccentricities, and their proximity to resonance. As explained in Section 3 of AD16, a comparison of TTVFaster with N-body integration shows better than 10% precision for a wide range of $\alpha-e$ parameter space given a planet pair. Although the model tends to fail near mean-motion resonances, for which $jP_{i}^{-1}\approx(j+k)P_{i+1}^{-1}$, the low eccentricities ($e_{1}\approx e_{2}\approx 0.01$) and mass-ratios ($\mu_{1}\approx\mu_{2}\approx 3\times 10^{-6}$) of our two planets indicate better than 1% precision for the observed planets [^3] Figures 5 and 6 of AD16 show the fractional precision of the analytic formulae compared to N-body integration. $\alpha\approx 0.723$ in our case.
Our pair-wise model allows us to decompose the transit times into the individual sources which perturb planets b and c. In Table 3, we list the peak TTV amplitudes ($A_{\text{TTV}}$) by perturbation source.
Table 3: Theoretical TTV peak amplitudes of solar system objects, calculated with TTVFaster.
Perturbation Source Affected Object $A_{TTV}$ [min]
EMB Venus 4.60
Mars Venus 0.19
Jupiter Venus 0.58
Saturn Venus 0.10
Venus EMB 3.44
Mars EMB 1.15
Jupiter EMB 3.21
Saturn EMB 0.38
2.2.3 Maximum likelihood
Now, we derive our maximum likelihood function for the fit of Eq. 4 to the transit times. In TTVFaster, a planet’s parameter set is composed of four orbital elements and one mass-ratio: $\left\{\mu_{i},P_{i},t_{0,i},e_{i},\omega_{i}\right\}$. We parameterize the eccentricity and the longitude of periastron using the Poincaré variables, $h\equiv e\sin{\omega}$ and $k\equiv e\cos{\omega}$, to create eccentricity vectors. The resulting parameter set for a system with $N_{p}$ planets, that are sorted by period, is given by $\textbf{x}=(\left\{\mu_{i},P_{i},t_{0,i},k_{i},h_{i}\right\};i=1,...,N_{p})$.
With respect to the initial conditions, we can rewrite the model times as $t_{\text{mod}}(\textbf{x})$, which we match with the ‘observed’ transit times from section 2.1.1 to compute the model likelihood. Let us define $\mathcal{L}(\textbf{x})$ as a function of the model parameters that represents the probability of measuring the observed data given the model, $p(\textbf{y}|\textbf{x})$. Each transit timing measurement has a Gaussian probability distribution and inaccuracies. Therefore, the log likelihood function for each data point is given by
$\log\mathcal{L}_{ij}(\textbf{x})=-\frac{\ln[2\pi(\sigma_{ij}^{2}+\sigma_{\text{sys}}^{2})]}{2}-\frac{(t_{\text{obs},ij}-t_{\text{mod},ij}(\textbf{x}))^{2}}{2(\sigma_{ij}^{2}+\sigma_{\text{sys}}^{2})},$ (5)
where $\sigma_{\text{sys}}$ is an unknown systematic uncertainty. By assuming that the timing errors are identical $(\sigma_{ij}=\sigma_{\text{obs}})$, we can rewrite the likelihood function in terms of the chi-square:
$\chi^{2}=\sum_{i=1}^{N_{p}}\sum_{j=1}^{T_{i}}\frac{(t_{\text{obs},ij}-t_{\text{mod},ij}(\textbf{x}))^{2}}{\sigma_{\text{obs}}^{2}}.$ (6)
Finally, the log likelihood for our dataset is
$\log\mathcal{L}(\textbf{x})=-\frac{N_{\text{obs}}\ln{[2\pi(\sigma_{\text{obs}}^{2}+\sigma_{\text{sys}}^{2})]}}{2}-\frac{\chi^{2}}{2}\frac{\sigma_{\text{obs}}^{2}}{\sigma_{\text{obs}}^{2}+\sigma_{\text{sys}}^{2}}.$ (7)
We can remove the likelihood’s dependency on the systematic uncertainty via marginalization of $p(\textbf{y}|\textbf{x})$. The result is expressed in terms of the incomplete gamma function; assuming that $\Gamma(N_{\text{obs}}/2-1)$ is much greater than $\Gamma(N_{\text{obs}}/2-1,\chi^{2}/2)$, the log likelihood function can be estimated by
$\log\mathcal{L}(\textbf{x})=\sum_{i=1}^{N_{p}}\sum_{j=1}^{T_{i}}\log\mathcal{L}_{ij}(\textbf{x})\approx\Big{(}1-\frac{N_{\text{obs}}}{2}\Big{)}\ln\chi^{2}\text{.}$ (8)
Next, we describe the models that are considered in this work.
Table 4: Summary of configurations used in our modeling of the simulated data. Each $P$ represents a planet around the stellar host.
Model $N_{p}$#\## Free Params.
$\mathcal{H}_{PP}$ 2 10
$\mathcal{H}_{PPP}$ 3 15
$\mathcal{H}_{PPPP}$ 4 20
2.2.4 Test configurations
For this paper, we test different planet configurations, and compute the maximum likelihood of various models given the transit times of Venus and the EMB. After finding the best 2-planet model parameters, we add trial objects to our model one at a time, in order to gauge where each hypothesis fails to describe the simulated transit times. We summarize each hypothesis in Table 4: $\mathcal{H}_{PPP}$ refers to our model with only three planets, and $\mathcal{H}_{PPPP}$ includes four planets. To illustrate, we show a diagram of our test configurations in Figure 2, all conditioned on two transiting planets.
A massive third planet (i.e. planet d) would impart the most gravitational force on the transiting planets. We add a planet with an initial mass ratio on the order of $10^{-3}$. We search 500 points in log-space for a peak in the period range 1826.25-8035.5 days, after which we refine the grid to better constrain planet d’s physical parameters.
A nearby terrestrial planet (i.e. planet e) would likely induce the next largest TTVs. We add a planet with a $10^{-7}$ stellar mass ratio, then search 500 points in log-space to find the most probable period within 584.4-1826.25 days. Once we find a peak in this wide range, we compute the likelihood around the peak period with 300 points in log-space. We use these computations to create the likelihood profile of a planet signal from the TTVs.
Figure 2: Schematic of the simulated multi-transit system, with the $\mathcal{H}_{PPP}$ and $\mathcal{H}_{PPPP}$ test configurations considered in this work. Aside from the observed planets (b and c), the large firebrick circle represents a gas giant, while the orange circle represents a terrestrial. Our search regions for additional planets are represented by the brackets.
Figure 3: Model comparison for simulated observations of Venus (planet b) and EMB (planet c): $\mathcal{H}_{PP}$ in dashed blue, $\mathcal{H}_{PPP}$ in solid orange, $\mathcal{H}_{PPP}$ in dot-dashed green. $\mathcal{H}_{PP}$ is insufficient for describing the transit times, especially for planet c, since the systematic uncertainty is more than 2 times the added noise level, $\sigma_{\text{obs}}$. The two panels on the right compare the model residuals to the injected TTV scatter. As shown, all models are consistent with a Gaussian (in dashed red) for planet b, but this is not so for planet c (i.e. EMB).
2.3 Error analysis
After completing a search for the period at maximum likelihood, we optimize our model with respect to the entire parameter set for each test configuration. This global fit provides the initial conditions for the MCMC sampling of our posterior probability. Mathematically, Bayes’ theorem states that $p(\textbf{x}|\textbf{y})=p(\textbf{y}|\textbf{x})p(\textbf{x})/p(\textbf{y})$ (Bayes & Price, 1763). In terms of the likelihood, we can rewrite this equation to give the posterior probability distribution, $P(\textbf{x})\propto\Pi(\textbf{x})\times\mathcal{L}(\textbf{x})$ where $\Pi(\textbf{x})$ is the prior function. While our approach to the likelihood in Eq. 8 does not account for the prior probability distribution, it provides an initial estimate with which to sample the posterior probability.
Excluding the eccentricity vectors, we place a uniform prior on the planetary parameters. High-multiplicity planetary systems have lower eccentricities in order to remain stable (He et al., 2020). Therefore, we place an informative prior that each planet in our model has an eccentricity between 0 and $e_{\text{max}}=0.3$, with a gradual decrease in probability from $e=0.2$ to $e_{\text{max}}$. To ensure a posterior that has a uniform probability distribution with eccentricity, we use $e_{i}=\sqrt{(e_{i}\cos{\omega_{i}})^{2}+(e_{i}\sin{\omega_{i}})^{2}}$ and a $1/e_{i}$ prior.
Using an affine-invariant Markov chain algorithm (Goodman & Weare, 2010), we sample our fitted parameter set and the systematic uncertainty, $\sigma_{\text{sys}}$. This parameter absorbs any inaccuracies in both our model and data analysis, resulting in a probability $P(\textbf{x},\sigma_{\text{sys}})$. Given that high dimensional Markov chains require a large number of steps, we use steps on the order of $10^{4}$ with 75 random walkers. After the burn-in – which we define as the first time the walkers cross the median parameter (Knutson, 2009) – the chain is examined to assess convergence and calculate the minimum effective sample size.
3 Results
In the following sections, we compare how well the test configurations from § 2.2.4 describe the transits of Venus and the EMB – henceforth dubbed planet b and c – which are perturbed by the remaining unseen SS bodies. Then, we explore the distributions of the retrieved masses and dynamics, and the effect of timing precision (§ 3.3).
Figure 4: Simulated TTVs of 2 transiting terrestrial planets – Venus/b (top) and the EMB/c (bottom) – with a best-fit linear ephemeris removed, colored by source. Left panels: The grey curve is the sum of all the perturbations, calculated from the mean values of the posterior sampling of the best-fit 3-planet model. Right panels: Same as the left, but here the solid grey line was created from the mean posteriors in our best-fit 4-planet model.
Figure 5: Upper panel: Likelihood profiles for Mars-like (orange curve) and Jupiter-like (firebrick curve) planets fit to the transit times of two terrestrial planets. The times used here were generated over 30 yrs with 30-sec noise. Lower panels: Comparison of likelihood (curve) and posterior probability (histogram) of orbital periods near the measured values for Mars (left) and Jupiter (right). We report the median systematic uncertainty, $\sigma_{sys}$, from the posterior sampling of each configuration, indicating that the four-planet model gives a better fit.
3.1 Example simulations and fits
In each of the four cases below, we use 80 transit epochs of Venus and EMB over 30 yrs, after injecting $\sigma_{\text{obs}}=30$-second noise. We plot the $O-C$ residuals from fitting this set of simulations in Fig. 3. Here, we see that the two planet model (in dashed blue) has very high residuals – with standard deviations of 40.0- and 123.9- seconds for planets b and c, respectively. That the model with only 2 planets is insufficient for describing the observed transit times over long time spans suggests that there must be additional planets. Indeed, the best-fit $\mathcal{H}_{PPP}$ model yields residuals with 35.9-s deviation for both planets, only 7.7 seconds higher than the actual standard deviation of the injected noise. Meanwhile, the $\mathcal{H}_{PPPP}$ residuals for each transiting planet have standard deviations of 35.9- and 23.7-s. In terms of error distribution, the models are generally consistent with each other and the Gaussian injected noise (see scatter panels of Fig. 3), barring the notably large residuals on $\mathcal{H}_{PP}$ for planet c.
We plot each source for the best-fit 3- and 4-planet configurations in Fig. 4, with different colors and line styles. These figures demonstrate the approximately 12 year periodic effects of Jupiter (firebrick curve), which are higher amplitude for planet c than on planet b. The grey lines represent the best analytic model fit to the observed times. The panels of Fig. 4 show the same simulated times, but with 3- or 4-planet fits; the dashed orange curve in the right panels shows the low amplitude TTV contribution of Mars. Interestingly, the pairwise TTVs of planets b and c are anti-correlated, demonstrating that they dynamically interact – this aids in breaking the model degeneracies that would be present if only one of these planets were seen to transit (Steffen & Lissauer, 2018).
3.2 The search for additional planets
Figure 5 shows the likelihood as a function of each orbital period for two non-transiting planets (i.e. a gas-giant and a terrestrial, aka Jupiter and Mars). The top panel of this plot displays where the peak probabilities occur based on a fit to the TTVs. The firebrick curves indicate the $\mathcal{H}_{PPP}$ model, while the orange curves are for the $\mathcal{H}_{PPPP}$. In the lower panels, we limit the x-axis to a narrow range near the true periods of Mars and Jupiter. These show a good agreement between the likelihood profile (curve) and the posterior histogram from MCMC (histogram) for planet 3 (Jupiter). The likelihood profile for the period of planet 4 has multiple peaks when coarsely fitting over the wide range – we discuss this result below.
For each assumption, we calculate the median posterior parameters within a 68.3% quantile interval, and the Bayesian Inference Criterion (BIC; Schwarz, 1978; Kass & Raftery, 1995). With respect to the maximum probability $P(\textbf{x},\sigma_{\text{sys}})$,
$\text{BIC}=-2\ln{[P_{\text{max}}(\textbf{x},\sigma_{\text{sys}})]}+k\ln{N_{\text{obs}}}\text{,}$ (9)
where $k$ is the number of free parameters in our sampling. Additionally, we compute the ${\chi}^{2}$ at the model parameters with the maximum probability – making sure to add the systematic uncertainty in quadrature. We compare our MCMC results with the true parameters for relevant SS objects in Table 5.
3.2.1 Two planets
The $\mathcal{H}_{PP}$ model has a BIC of $=-1013.54$. Although the systematic uncertainty is 82.1 seconds, the physical parameters are accurate to those of Venus and Earth: $m_{b}=0.836^{+0.151}_{-0.136}$ and $m_{c}=1.002^{+0.106}_{-0.100}$. Additionally, we derive eccentricities that are consistent with the actual values: $e_{b}=0.005_{-0.011}^{+0.016}$ and $e_{c}=0.009_{-0.007}^{+0.0011}$.
Figure 7 displays a corner plot of the mass-ratios, periods, and eccentricity-vectors of this configuration. The periods and masses follow Gaussian distributions, while the constraints on the eccentricities for the two planets are strongly correlated. We attribute this aspect to the eccentricity-eccentricity correlation described in Lithwick et al. (2012).
3.2.2 Three planets
[…]
We find the same relations for the masses and eccentricities of the transiting planets as in the $\mathcal{H}_{PP}$ model (Figure 8). In addition, $e\cos{\omega}$ for planet d (i.e. Jupiter) has a negatively skewed tail. When this component is plotted against the mass-ratio, we find a non-linear correlation. This approximate degeneracy between masses and orbit shapes is known to prevent measuring planet masses due to observations with insufficient timing precision (see Hadden & Lithwick, 2017; Leleu et al., 2023).
Despite this degeneracy, the addition of this third planet helps to better constrain the whole system thanks to the lower systematic uncertainty of $\sigma_{sys}\approx 25$ sec. Compared to the 2-planet configuration, the masses and eccentricities of planets b and c are even better constrained: $m_{b}=0.840^{+0.064}_{-0.062}$, $e_{b}=0.001_{-0.004}^{+0.007}$; and $m_{c}=1.007^{+0.044}_{-0.043}$, $e_{c}=0.014_{-0.003}^{+0.004}$.
3.2.3 Four planets
Recall that we found multiple peaks in the likelihood profile, with the highest probability at $687.2$ days (Fig. 5, upper panel). The lower left panel of Fig. 5 reveals that the posterior period distribution closely agrees with the likelihood profile near the peak. However, the traces for $P_{e}$ suggested that the 4-planet model had not sufficiently converged. We ran the chains with 20,000 additional steps, but found that many walkers continue to sample low probability regions with large systematic uncertainties.
[…]
For completeness, we plot the decomposed TTVs for this fit in Fig. 4, and report the planet masses in Tab. 5: $m_{b}=0.830_{-0.058}^{+0.059}$ and $m_{c}=1.002\pm{0.039}$, $m_{d}=284.345^{+55.829}_{-52.365}$, and $m_{e}=0.093^{+0.049}_{-0.040}M_{\oplus}$. Notably, our measured mass for planet d becomes more accurate with that known for Jupiter, albeit with larger uncertainties. The large uncertainty on planet e’s mass is driven by the broad tails of the posterior probability which are far from the correct period of Mars, and thus are not retrieving its correct mass. Therefore, we conclude that the analytic model is simply not capable of constraining the errors on the properties of planet e (i.e. Mars) for this noise level and observing span.
3.2.4 Five planets
Due to its large distance from the transiting planets, a Saturn-like planet would only induce TTVs on the order of 10s of seconds on Venus and Earth (see Table 3). As this is reaching our most optimistic observing limits, we do not include its retrieval in this work. We did perform an exploratory fit with five planets, and found the following: between the orbital periods of 26 and 40 years, the maximum likelihood peaks at 32 years. We report the following mass-ratios for this fit: $\mu_{b}\approx 2.3795\times 10^{-6}$, $\mu_{c}\approx 3.0474\times 10^{-6}$, $\mu_{e}\approx 0.2597\times 10^{-6}$, $\mu_{d}\approx 0.0010$, $\mu_{f}\approx 0.0009$. The orbital periods for the unseen planets are $P_{d}=4327.890$, $P_{e}=688.277$, $P_{f}=9429.632$ in days, while the eccentricities are $e_{d}=0.052$, $e_{e}=0.155$, $e_{f}=0.092$.
Table 5: Retrieved parameters from TTVs in Table 2, listed by test configuration. For each posterior parameter, we give the median value after burn-in and the $68.3\%$ confidence region. In the right-most column, we provide masses and Keplerian orbital elements for Venus, EMB, Jupiter and Mars – adopted from Hussmann et al. (2009) or calculated using JPL ephemerides (Park et al., 2021). Time of periastron passage, $t_{0}$, is expressed as $\text{JD}-2430000$. The BICs are calculated by finding the maximum likelihood of the posterior sample.
Parameter $\mathcal{H}_{PP}$ $\mathcal{H}_{PPP}$ $\mathcal{H}_{PPPP}$ Truth
BIC -1013.54 -1130.33 -1138.12
$\chi^{2}$ 73.44 74.174 77.519
Degrees of Freedom 70 65 60
Fitted Parameters
$\sigma_{sys}$ [days] $0.000945_{-8.530\times 10^{-5}}^{+9.654\times 10^{-5}}$ $0.000285_{-6.523\times 10^{-5}}^{+6.518\times 10^{-5}}$ $0.000189_{-9.649\times 10^{-5}}^{+8.466\times 10^{-5}}$
Planet b
$\mu_{b}\times 10^{-6}$ $2.5102_{-0.4098}^{+0.4537}$ $2.5243_{-0.1875}^{+0.1926}$ $2.4927_{-0.1749}^{+0.1765}$ $2.4464$
$P_{b}$ [days] $224.700778_{-0.000009}^{+0.000010}$ $224.700780_{-0.000005}^{+0.000005}$ $224.700778_{-0.000005}^{+0.000005}$ $224.7007$
$t_{0,b}$ [days] $3503.765444_{-0.000287}^{+0.000275}$ $3503.765354_{-0.000129}^{+0.000128}$ $3503.765396_{-0.000131}^{+0.000124}$ $3503.7644$
$e_{b}\cos{\omega_{b}}$ $-0.0047_{-0.0171}^{+0.0190}$ $-0.0009_{-0.0075}^{+0.0056}$ $-0.0013_{-0.0066}^{+0.0047}$ $-0.003$
$e_{b}\sin{\omega_{b}}$ $0.0003_{-0.0111}^{+0.0157}$ $0.0006_{-0.0049}^{+0.0074}$ $0.0012_{-0.0042}^{+0.0068}$ $-0.006$
Planet c
$\mu_{c}\times 10^{-6}$ $3.0112_{-0.2991}^{+0.3173}$ $3.0248_{-0.1304}^{+0.1320}$ $3.0094_{-0.1160}^{+0.1185}$ $3.0369$
$P_{c}$ [days] $365.256433_{-0.000021}^{+0.000020}$ $365.256454_{-0.000011}^{+0.000010}$ $365.256455_{-0.000010}^{+0.000010}$ $365.256355$
$t_{0,c}$ [days] $3624.402555_{-0.000361}^{+0.000360}$ $3624.402177_{-0.000176}^{+0.000178}$ $3624.402149_{-0.000164}^{+0.000166}$ $3624.4054$
$e_{c}\cos{\omega_{c}}$ $0.0092_{-0.0124}^{+0.0162}$ $0.0131_{-0.0058}^{+0.0047}$ $0.0126_{-0.0051}^{+0.0040}$ $0.011$
$e_{c}\sin{\omega_{c}}$ $0.0017_{-0.0086}^{+0.0126}$ $0.0032_{-0.0042}^{+0.0057}$ $0.0040_{-0.0036}^{+0.0052}$ $0.012$
Planet d
$\mu_{d}$ $0.000711_{-0.000138}^{+0.000140}$ $0.000854_{-0.000157}^{+0.000168}$ $0.00095$
$P_{d}$ [days] $4206.06_{-58.80}^{+58.35}$ $4328.83_{-67.23}^{+58.89}$ $4332.82$
$t_{0,d}$ [days] $677.866_{-155.363}^{+159.057}$ $389.425_{-147.851}^{+174.062}$ $333.7268$
$e_{d}\cos{\omega_{d}}$ $0.0022_{-0.0295}^{+0.0227}$ $0.0093_{-0.0239}^{+0.0193}$ $0.0403$
$e_{d}\sin{\omega_{d}}$ $-0.0351_{-0.0196}^{+0.0174}$ $-0.0403_{-0.0155}^{+0.0138}$ $0.0268$
Planet e
$\mu_{e}\times 10^{-6}$ $0.2794_{-0.1204}^{+0.1475}$ $0.3227$
$P_{e}$ [days] $687.373_{-2.774}^{+3.032}$ $686.980$
$t_{0,e}$ [days] $401.504_{-43.106}^{+36.181}$ $383.823$
$e_{e}\cos{\omega_{e}}$ $-0.0584_{-0.0519}^{+0.0932}$ $0.0131$
$e_{e}\sin{\omega_{e}}$ $-0.0857_{-0.0669}^{+0.0587}$ $0.0925$
Derived Parameters
$m_{b}[M_{\oplus}]$ $0.836^{+0.151}_{-0.136}$ $0.840^{+0.064}_{-0.062}$ $0.830^{+0.059}_{-0.058}$ $0.815$
$m_{c}[M_{\oplus}]$ $1.002^{+0.106}_{-0.100}$ $1.007^{+0.044}_{-0.043}$ $1.002^{+0.039}_{-0.039}$ $1.012$
$m_{d}[M_{\oplus}]$ $236.563^{+46.680}_{-45.810}$ $284.345^{+55.829}_{-52.365}$ $317.8$
$m_{e}[M_{\oplus}]$ $0.093^{+0.049}_{-0.040}$ $0.107$
$e_{b}$ $0.005_{-0.011}^{+0.015}$ $0.001_{-0.004}^{+0.007}$ $0.002_{-0.004}^{+0.007}$ $0.006$
$e_{c}$ $0.009_{-0.007}^{+0.011}$ $0.014_{-0.003}^{+0.004}$ $0.013_{-0.004}^{+0.004}$ $0.016$
$e_{d}$ $0.035_{-0.020}^{+0.018}$ $0.041_{-0.021}^{+0.018}$ $0.048$
$e_{e}$ $0.104_{-0.078}^{+0.067}$ $0.093$
  • Notes. Murray & Dermott (2000) provides equations for calculating solar system orbital elements at a given JD. However, we do not employ these in our calculation of the true eccentricity vectors since TTVFaster computes instantaneous parameters while the provided equations utilize secular rates of change per century. Instead, we calculate the $\omega$ components of the true eccentricity vectors by finding the angular difference in longitude for each planet between J2000 and the “observed” time of periastron passage – which is defined as when both transiting planets had similar mean longitudes ($\lambda_{1}\sim\lambda_{2}$). We report this angular difference as 77.
Figure 6: The effect of timing precision on the retrieved masses for the best-fit $\mathcal{H}_{PPP}$ (top) and $\mathcal{H}_{PPPP}$ (bottom) models. The true masses for Venus, EMB, Jupiter, and Mars are plotted as horizontal lines. The measured masses of planets b and c are well constrained when fitting three planets, even at the 60-s level. When fitting four planets, the mass of Jupiter is more accurate; but we are unable to probe noise levels greater than $\approx 35$ seconds. The x-axis shows the injected noise and the systematics found from the Markov chains.
Figure 7: Posterior probability density functions of 8 components of the $\mathcal{H}_{PP}$ model, which includes the mass-ratios, periods, and eccentricity-vectors of each planet. The contours enclose the regions that contain 0.683 and 0.955 of the cumulative posterior probability. In the panels along the diagonal, the green shaded regions bounded by dashed vertical lines correspond to the 68.3% confidence interval of a given component. Contours are based on 8261 sample steps.
Figure 8: Posterior probability density functions of 12 components of the $\mathcal{H}_{PPP}$ model, which includes the mass-ratios, periods, and eccentricity-vectors of each planet. Same notes as Fig 7, but the contours are based on 26857 sample steps.
Figure 9: BICs for $\mathcal{H}_{PP}$ (right) and $\mathcal{H}_{PPP}$ (left): darker bins have lower BICs, which indicates a more preferred solution. For the regions of overlap, the minimum change in BIC is 39, corresponding to a $6.3\sigma$ detection of a Jupiter-analogue. Empty bins indicate where we did not have convergent Markov chains.
Figure 10: Posterior probability density functions of 16 components of the $\mathcal{H}_{PPPP}$ model, which includes the mass-ratios, periods, and eccentricity-vectors of each planet. Same notes as Fig 7, but the contours are based on 24975 sample steps.
3.3 Masses and orbital elements
The simulations in the previous subsection are all for 30 year baselines with 30 seconds of Gaussian noise injected. This case was selected because it allowed for a direct comparison between the test configurations. In Table 5, we report the retrieved parameters for each configuration in comparison to the truth. As more planets are included, the uncertainties on variables fit in the previous model are improved. This is especially important for measurements of planetary masses and orbital shapes (i.e. eccentricities). We find that these derived parameters are accurate for planets b, c, and e; planet d’s mass is an underestimate until planet e (i.e. Mars) is included in the configuration.
In our search for additional planets – $\mathcal{H}_{PPP}$ and $\mathcal{H}_{PPPP}$ – we cannot report uncertainties due to a lack of chain convergence when the survey baseline is shorter than 22 years. This is also apparent with noise levels greater than 30 seconds (for $\mathcal{H}_{PPPP}$) and 90 seconds (for $\mathcal{H}_{PPP}$). Therefore, we could not directly compare our test configurations for the entire simulation set. We believed that given longer chains these runs could eventually converge to reasonable values. However, the analytic model may just be insufficient for such cases.
3.4 Effect of timing noise and observing baseline
For the $\mathcal{H}_{PPP}$ case with 30-year duration, we plot how timing precision affects our posterior masses in the top panel of Figure 6. Although an increase in injected noise leads to larger mass uncertainties for a given model, the measured masses for planets b and c remain consistent across all of the models. The lower panel shows how timing precision affects the masses in the $\mathcal{H}_{PPPP}$ case, where available.
When we examine our entire simulation set, as described in Section 2.1, we find that the observing baseline is not the dominant factor in determining the mass uncertainties at high timing precision. Figures 11 and 12 show how the percent error in the mass-ratio measurements for planets b, c, and d changes for a given simulated survey configuration. For the same simulations (i.e. same realizations of noise), the addition of the third planet to the model greatly reduces the percent errors for planets b and c. For the example case, there is a reduction from 11% to 4% error for the EMB after adding a third planet to our TTVFaster model. Likewise, we can reduce the 20% error on the measurement of Venus’ mass to 8%. In contrast to these two planets, the gas giant’s parameters are usually much less accurate and precise in the 3-planet model (see Fig. 12 and 6).
For the simulations that are directly comparable, we calculate the BIC and plot this value in Fig. 9. Not only is $\mathcal{H}_{PPP}$ always a better fit to the transit times, we find at least a 6-$\sigma$ detection of Jupiter with a minimum 22-yr baseline and a 60-s noise level.
[…]
Figure 11: Percent error in average mass-ratio measurements for planets b (left panel), and c (right panel), from modelling 2-planets. The 5 bins for the y-axis correspond to the injected noise levels: 10-, 30-, 60-, 90-, and 120-seconds. Empty bins indicate where we did not have convergent Markov chains. We use different colorbar scales for each parameter, and we only report results for the bins where the percent error was less than 100%.
Figure 12: Same as in Fig. 11, but here we plot the percent errors of the best 3-planet model: planets b (left panel), c (middle panel), and d (right panel). Compared to the model with only 2 planets, there are much lower percent errors in the mass-ratios for planets b and c.
Figure 13: $\Delta$BIC for $\mathcal{H}_{PPP}$ and $\mathcal{H}_{PPPP}$). Darker bins show more confidence in detecting Mars; high significance is only feasible when we inject 10-seconds of timing error.
Figure 14: Measured planet mass as a function of orbital period for multi-planetary systems with $N_{p}$ planets, listed by discovery method: transit (green circles), RVs (blue squares), TTVs (purple triangles), and others (yellow dots). The mass measurements in this work – for Venus, EMB, Mars, and Jupiter – are taken from the lower panel of Figure 6: the 4-planet model with a total uncertainty of $\approx$ seconds. Based on data from the NASA Exoplanet Archive accessed on 2024 May 29 (NASA Exoplanet Science Institute, 2020).
4 Discussion
4.1 Model Accuracy
In this work, we find that the retrieved posteriors of the transiting planets are consistent across the test configurations. Specifically, for Venus and the EMB, the measured masses stay consistent with the correct values across the models with 2-4 planets (Table 5). In addition, the uncertainties on the measured masses of the planets improve as more planets are included in the model. This is due to the fact that the 3- and 4-planet models better reproduce the actual TTVs, and this results in a smaller systematic error, hence improving the precision of their masses. In addition, as the timing precision improves, the error bars on the measured masses become more precise for the 3-planet model (Fig. 6). Thus, we expect that a long-term campaign to measure the masses of the pair of transiting planets is robust to the timing uncertainty and the number of planets in the model; that is, the inferred masses should be accurate. In addition, the precision of the masses should improve as the timing errors become more precise, and as the model becomes more precise with the inclusion of more planets. The same applies to these planets’ eccentricities, which become more precise and accurate as the model is improved from 2- to 3- to 4- planets.
As far as Jupiter is concerned, our results are mixed. In the 3- and 4- planet models, we detect Jupiter at high significance near the correct orbital period for the 30-second, 30-year survey, as well as surveys with shorter durations and larger timing noise levels (see Figure 12). However, the mass of Jupiter is not retrieved accurately in the 3-planet model until high timing precision is achieved (Figure 6), or when a 4-planet model is used (Table 5).
This indicates that the detection of a giant planet with TTVs is more straightforward than its characterization. Jupiter’s eccentricity is retrieved more accurately and precisely in the 4-planet model than in the 3-planet model, and its orbital period and transit time (or time of inferior conjunction if it does not transit) are consistent with the correct values in the 4-planet case, but not in the 3-planet case. This indicates that despite its small mass, the presence of Mars does have an effect on the inference of the correct parameters of Jupiter; Mars is small, but it is (relatively) mighty. This can be attributed to the closer proximity of Mars to the Earth-Moon-Barycenter, as well as its proximity to a 2:1 resonance, both of which enhance its dynamical influence on the EMB relative that of Jupiter, whose long orbital period greatly diminishes its dynamical impact on Earth. Nevertheless, Mars is 3.5 orders of magnitude smaller in mass than Jupiter, so after Venus, Jupiter still dominates the TTVs of the EMB, but Mars’ influence does interfere in the 3-planet analysis.
4.2 Implications for observational capabilities
To find systems with architectures analogous to our solar system, we need to improve the transit timing sensitivity of a telescope to better than the 30-s noise level for Earth/Venus/Sun-analog transits. If we can could achieve this, we could probe a region of exoplanet discovery space which is unreachable with past and current technology. Figure 14 shows the measured masses for known planets by discovery method. We include masses derived in this work for a 30 year observing span with 10-s injected noise fit with the $\mathcal{H}_{PPPP}$ model in red – plotted on top of the true SS values. Compared to other multi-planet systems, this work provides the only way to characterize a system with both terrestrial and gas giant planets.
Unfortunately, there are no existing instruments that can detect and measure the TTVs over the baselines required to detect a SS analog. However, future telescopes may have this capability. Matuszewski et al. (2023) have estimated that about a dozen Earth-sized planets with orbits between 250-500d may be possible to find Earth-Sun twins with PLATO or Nautilus (Apai et al., 2019). Additionally, Borsato et al. (2022) claim that the upcoming Ariel mission can obtain the required timing precision ($\simeq 30$-s) for sufficiently bright stars – such as 55 Cnc with $V=5.95$ mag, and K2-24 with $V=11.2$ mag – using its fine guidance sensors.
With a transit time precision at (or better than) 60 sec, our analytic model still strongly detects a planet beyond 3 au, via its perturbations of the inner transiting planets. This suggests that we do not require a fortuitous microlensing signal in order to discover distant planets like those in our SS (Poleski et al., 2014) . In any case, a space mission would still need to be long-lived in order to properly analyze the long term TTVs – at least one to two orbits of the wide-separation planet.
Radial velocity monitoring of a multi-transiting host star would improve mass constraints for the outer gas giant (i.e. Jupiter in our example system). In the case of an extra-solar system, one could measure the RV signals caused by the non-transiting planet, if RV data are sufficiently precise (see Pepe & Lovis, 2008; Pepe et al., 2013), which could complement and confirm a transit-timing analysis, and vice versa. One potential candidate for such a study is TOI-700 with four transiting planets in the TESS continuous viewing zone, albeit orbiting a lower mass M dwarf (Gilbert et al., 2020, 2023). This system has two terrestrial-sized habitable-zone planets in close proximity to one another (Gilbert et al., 2023). Intense radial-velocity and transit-timing follow-up of this system might enable the detection of a more distant planets with both techniques, as well as the mass characterization of the transiting planets.
For context, He et al. (2021) report that extensive RV monitoring over several years would be required to measure the masses of Venus and Earth analogs: 200 RV observations needed to measure the semi-amplitude (and therefore the mass) of a transiting Venus in a Kepler-like system with better than 20% precision using a next-generation RV instrument ($10$ cm/sec precision). This level of precision is on par with what we retrieve for a transiting Venus, for transit times with $\approx 30$ sec errors). As shown in Figure 14, using TTVs of an MTS presents the only current way to accurately probe unseen long period terrestrials. These planets, although small, create effects that can be constrained by sampling the TTV super-period – as Mars does on Venus and the Earth-Moon-Barycenter.
4.3 Future work
In its current state, TTVFaster breaks down in accuracy at high eccentricities, for large planet-star mass ratios, and for planets close to resonance (Agol & Deck, 2016). Therefore, future work could involve implementing a more accurate model, such as the NBodyGradient code (see Agol et al., 2021) to analyze the simulated solar system. Similarly, a full photo-dynamical simulation of realistic transits with varying levels of white noise and correlated noise appropriate for our Sun could help us determine how precisely the times can be measured (a la Morris et al., 2020, but with bigger glass or multiple wavelengths to achieve higher precision). Adding in the Moon to our analysis is another necessary next step. Investigating the impact of missing transits and transit-timing outliers we also leave to future work.
5 Conclusion
To date we have yet to detect and characterize an exoplanet system that is analogous to our Solar System. In principle it should be possible to detect the transits of both Earth and Venus as seen from afar, and so here we have investigated what might be learned about their masses and the architecture of the Solar System from long-term, high-precision transit-timing.
In this study, we used an approximate analytic model for transit timing variations (AD16) to retrieve the properties of planets in a simulation of the Solar System with two detected transiting planets: Venus and Earth-Moon-barycenter (we leave the influence of the Moon on Earth to future work). The AD16 model is plane-parallel, which is suitable for our analysis because mutual inclinations and eccentricities of the relevant Solar System bodies are small.
Our focus was to compute the simultaneous detectability of an additional non-transiting gas giant and terrestrial planet, given detectable TTV signals from two transiting rocky planets. The results are as follows:
  • The TTVs for EMB and Venus are anti-correlated. This anti-correlation enables the detection of additional planets which affect each of the transiting planets differently, overcoming degeneracies which are present when only one transiting planet displays TTVs (Lithwick et al., 2012).
  • We recover the correct masses and orbit shapes for Venus and EMB with less than 50 observations, each. Their measured masses appear to be robust; that is, their masses are accurate whether or not we include additional planets in the transit-timing model. However, their masses become more precise when we include Jupiter in the transit-timing model. This indicates that perturbations by an unknown planet can lead to an additional source of uncertainty in transit-timing models.
  • We can readily infer the existence of a wide-separation gas giant (i.e. Jupiter), with up to 90 sec transit timing noise, with a duration of only 1.25 orbits of Jupiter observed.
  • The transit timing noise level required to measure a reliable mass ($<30\%$ error) for Jupiter is 60 seconds.
  • The derived orbital period and mass for Jupiter are underestimates of the actual values, unless we have very precise timing (10-sec), really long baselines ($>26$ yrs), and/or include Mars in our transit-timing model.
  • With the addition of Jupiter to our model, we can measure the masses of Venus and EMB with uncertainties of 8% and 4%, respectively, for a 30-year survey with 30-second timing precision. But, even when analyzing a model only accounting for two planets, or analyzing simulations with larger timing uncertainties, we still obtain accurate masses for Venus and EMB, albeit with larger uncertainties.
  • At very high timing precision ($\sigma_{\text{obs}}=10$), we can detect Mars at the correct period.
  • We would require better than 30 sec timing precision to detect and characterize a system analogous to ours – but only for the planets out to several astronomical units, excluding Mercury (Figure 14).
Acknowledgements
B.L. is supported by the NSF Graduate Research Fellowship, grant no. DGE 1762114. E.A. acknowledges support from NSF grant AST-1907342, NASA NExSS grant No. 80NSSC18K0829, and NASA XRP grant 80NSSC21K1111.
This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.
Appendix A Solar System Motions
A.1 Required Observer Location
In order to determine transit times, we must calculate a multi-transit vector, which points to the location from which a distant observer could detect transits of multiple planets in front of their host star. To do so, we must first find the location from which one would see two planets eclipsing a star. Let the position vector $\textbf{r}=(x,y,z)$ specify the distance $r$ and direction $\hat{\textbf{r}}$ of a planet from a star. For the $i$th planet at position $\textbf{r}_{i}$, the velocity and acceleration vectors are defined as follows: $\textbf{v}_{i}\equiv\dot{\textbf{r}}_{i}$, and $\textbf{a}_{i}\equiv\ddot{\textbf{r}}_{i}$. We define $\textbf{h}_{i}$ as the angular momentum integral: the vector normal to the orbital plane containing the $\textbf{r}_{i}$ and $\textbf{v}_{i}$ vectors. In a two-planet system, the unit vector normal to the angular momentum integrals of both planets is
$\hat{\textbf{n}}_{\text{obs}}=\pm\frac{\textbf{h}_{2}\times\textbf{h}_{1}}{|\textbf{h}_{2}\times\textbf{h}_{1}|}=\frac{(\textbf{r}_{2}\times\textbf{v}_{2})\times(\textbf{r}_{1}\times\textbf{v}_{1})}{|(\textbf{r}_{2}\times\textbf{v}_{2})\times(\textbf{r}_{1}\times\textbf{v}_{1})|},$ (A1)
and points in the direction where an observer could see both Venus ($i=1$) and the Earth-Moon Barycenter ($i=2$). We illustrate this vector in Figure 15, where the orbital planes of each object intersect.
A.2 Transit Definition
A planet would be seen to transit its host star when its projected distance to the star on the sky plane is smallest, it is in front, and that minimum separation is less than the sum of the radii of the two bodies. At this instance, the planet’s sky velocity is perpendicular to its sky position such that the vectors $\textbf{r}_{\text{sky,}i}$ and $\textbf{v}_{\text{sky,}i}$ are orthogonal. Minimizing the projected planet-star distance amounts to solving
$f\equiv\textbf{r}_{\text{sky,}i}\cdot\textbf{v}_{\text{sky,}i}=0$ (A2)
at the conjunction. Note that this transit condition also applies when the planet is occulted by the star, unseen by the observer. We ensure that the observer sees the planet in front of the star during the transit by imposing two additional conditions: 1) the first derivative $f^{\prime}>0$ and 2) $\textbf{r}_{i}\cdot\hat{\textbf{n}}_{\text{obs}}>0$.
The sky plane is perpendicular to the observer’s line of sight. Given an observer in the direction $\hat{\textbf{n}}_{\text{obs}}$, we can compute when this observer will detect the mid-point of transit by rewriting the function $f$ at a time $t$. Therefore, the transit condition becomes
$f(t)=(\textbf{r}_{i}\cdot\textbf{v}_{i})-(\textbf{r}_{i}\cdot\hat{\textbf{n}}_{\text{obs}})(\textbf{v}_{i}\cdot\hat{\textbf{n}}_{\text{obs}}),$ (A3)
and the derivative is
$f^{\prime}(t)=(\textbf{v}_{i}\cdot\textbf{v}_{i})+(\textbf{r}_{i}\cdot\textbf{a}_{i})-(\textbf{v}_{i}\cdot\hat{\textbf{n}}_{\text{obs}})^{2}-(\textbf{r}_{i}\cdot\hat{\textbf{n}}_{\text{obs}})(\textbf{a}_{i}\cdot\hat{\textbf{n}}_{\text{obs}}).$ (A4)
If $f^{\prime}(t)$ changes sign from negative to positive during a time step of the orbital integration while the above requirements are met, a transit has occurred (Deck et al., 2014).
Figure 15: Schematic of a Sun-Venus-EMB system: the orbital planes of Venus and the Earth-Moon Barycenter are represented by ellipses; and the misalignment of these orbits is exaggerated for demonstration purposes. The mean angular difference between the inclinations of these orbits is 3.394 at the epoch of J2000 (i.e. JD 2451545.0).